OpenStudy (anonymous):

At a jewelry store, the price P of a gold necklace varies directly with its length l. Also the weight w of a necklace varies directly with its length. Show that the price of a necklace varies directly with its weight.

OpenStudy (anonymous):

\[ P=kl \]And \[ w = cl \]so \[ \frac wc=l \]And \[ P = k\frac wc=\frac kc w \]

OpenStudy (anonymous):

So if you let \(k' = \frac kc\)\[ P=k'w \]

OpenStudy (anonymous):

Which is direct variation.